a car is travelling at 30m/s and has kinetic energy of 450KJ, what is it's mass?


1
Expert's answer
2021-03-08T08:30:45-0500
"KE=\dfrac{1}{2}mv^2,""m=\dfrac{2KE}{v^2}=\dfrac{2\cdot450\cdot10^3\ J}{(30\ \dfrac{m}{s})^2}=1000\ kg."